OpenStudy (anonymous):

water has 5 times the heat capacity of soil

OpenStudy (anonymous):

great!

OpenStudy (unklerhaukus):

\[{c_p}_{H_2O} = 4.18 \left[\frac{\text{J}}{g·K}\right]\] \[{c_p}_{H_2O}=5\times {c_p}_{\text{soil}}\] \[{c_p}_{\text{soil}}=\frac{{c_p}_{H_2O}}5=\frac{4.18}{5}\left[\frac{\text{J}}{g·K}\right]=0.836\left[\frac{\text{J}}{g·K}\right]\]